OpenStudy (anonymous):

A plane travels from Orlando to Denver and back again. On the five-hour trip from Orlando to Denver, the plane has a tailwind of 40 miles per hour. On the return trip from Denver to Orlando, the plane faces a headwind of 40 miles per hour. This trip takes six hours. What is the speed of the airplane in still air?

OpenStudy (hba):

this is how i would solve it it is distance equals rate x time a is the speed in still air the rate in the tailwind is a+40 the rate in the headwind is a-40 time is 5 hours in tailwind time is 6 hours in headwind 5 x (a+40)= 6 x (a-40) solve and you should get a= 360mph

OpenStudy (ganpat):

Tailwind is wind with object, while Headwind is wind against the direction so, 5 (s + 40) = 6 (s - 40) 5s + 200 = 6s - 240 s = 440 mph...